### v4

@@ 4.1.2 Supervised ML Approaches
The randomisation must decrease correlation while retaining strength to increase accuracy. The forests under study grow each tree by applying inputs or combinations randomly chosen at each node. Accuracy results from the produced forests are favourably comparable to those of Adaboost. This class of methods has the following desirable qualities:
  • Its accuracy is comparable to Adaboost and occasionally much better.
  • It is relatively resistant to noise and outliers.
  • It moves more quickly than bagging or boosting.
  • It is straightforward, easily paralleled and provides helpful core estimates of error, strength, correlation, and variable relevance.
@@
LR is a linear model used to classify instead of predict. A logistic function is used for probabilities that describe the possible results of a single trial in this model. Typically, a sigmoid curve with the shape of a Sigmoid curve is called a logistic function or logistic curve. Wikipedia (2023)
XGBoost:
XGBoost is built using decision trees that have increased their gradients. Compared to other ML algorithms, XGBoost is fast and performs well. It is a technique for boosting machinery or doing machine boosting. Extreme Gradient Boosting, or XGBoost, helps maximise memory and hardware resources when used with tree-boosting techniques. The algorithm is enhanced, and the model is changed, which has benefits. The second derivative is used to perform Taylor expansion on the cost function, improving the results’ accuracy. The XGBoost method employs an additive training approach to optimise the objective function, suggesting that the outcome of the earlier stage affects the optimisation process in the subsequent stage. Gad et al. (2021)
@@ 4.2 Technical Implementation Methodology
This study will address multiclass problems using two approaches: Reducing Multiclass to Binary decision and multiclass methodology. The performances of the two approaches will then be discussed below.
Binary decision reduction:
Before feeding the data into ML algorithms, cleaning and preparation are crucial steps to achieve high performance. Unnecessary features may impact the performance of the chosen ML method. Missing values are regularly present in datasets.
To provide a relevant analysis, these values must be handled properly. After grouping the same attacks and splitting the datasets into two parts, training data and testing data, a Binary classifier was used to label the dependent columns in BENIGN and ATTACK. LabelEncoder is used to normalise labels so that the dataset only contains values between 0 and n classes-1. The model fitting and learning purposes are not similarly affected by variables measured at different scales and might lead to bias. Consequently, feature-wise normalisation, like the MinMax Scaling, is typically performed before model fitting to address this potential issue. Normalizer is applied row-by-row, while MinMaxScaler is done column-by-column. It balances all the information structures in the range [0, 1], otherwise in the range [-1, 1] in case there are negative values in the database.
General Multiclass methodology:
@@
To begin this procedure, distinct datasets were constructed by keeping a specific attack and dropping the remainder of the attacks to create a new data combination for each particular attack vs BENIGN data.
Feature Selection:
@@
The SelectKBest class from the scikit-learn library uses various statistical trials to select specific features. For example, the experiment used the chi-squared (chi2) statistical test as a scoring function for non-negative characteristics to pick the ten best features. The Chi-Square test is a statistical method for measuring the discrepancy between observed and projected data. This test can also be achieved to see if the data’s categorical variables are related. For example, it is helpful to determine whether a discrepancy between two categorical variables results from coincidence or an association between them. The Chi-Square Test is a method for quantifying the degree of disparity between expected and actual findings. To determine whether the results that were observed and those that were anticipated differed, the Chi-Squared statistic was used.
Hyperparameter Optimization: Bayesian Optimization with Gaussian Process:
BayesSearchCV incorporates "fit" and "scoring" methods. Using a cross-validated search of parameter settings, the parameters of the estimator used to apply these methods are optimised. In contrast to GridSearchCV, several parameter settings have been experimented with from the selected distributions rather than all parameter values. n_iter requires the number of parameter settings that are assessed. The ideal hyperparameters for the models can be found by efficiently scanning the hyperparameter space by utilising optimisation techniques throughout the design process of ML models. An estimator (a regressor or classifier) with its optimisation problem, a search space (configuration space), a search or enhancement method used to find hyperparameter combinations, and an evaluation function used to evaluate the performance of various hyperparameter setups. A hyper-domain parameter can be categorical, binary, discrete, continuous, or discrete. As a result, there are three types of hyper-parameters: continuous, discrete, and categorical. In real-world applications, continuous and discrete hyperparameters’ domains are typically constrained. On the other hand, conditionality can occasionally be seen in the hyper-parameter configuration space. Depending on the value, using or tuning a hyperparameter may be necessary. This study used hyperparameter optimisation to feed stacking base classifier RF.
Ensemble Stacking Classifier:
There are three different methods of ensemble learning: bagging, boosting, and stacking. The alternate voting methods are bagging and boosting. Standardised models are used in the bagging strategy to estimate the test data class. Predictions from homogeneity chosen models are initially recorded. The test data are then assigned to the class that most models predict. Similar to boosting, the training phase of the models includes extensive training for the misclassified data. The model with the highest accuracy is finally chosen as the classifier for the test data. Because stacking is a generic framework with numerous ensemble approaches, it is a successful strategy. Base and meta-learning are two of the learning layers. A training data set is used to teach the initial (base) learners in base learning. After completion of training, the base learners generate a new dataset for the meta-learner. The meta-learner is then trained using a clean set of training data. Finally, the test set is categorised using the trained meta-learner. The ideal foundation learner should be chosen as it is an essential stacking component. In other words, numerous base learners should be used for the training data set rather than just one. The primary distinction between stacked and other ensemble approach methods is the application of meta-level learning-based classification as the final classification in stacking. As mentioned, this study utilised random forest (RF) and XGboost as base learners in a stacked ensemble model. Lastly, the final learner is logistic regression.
Experimantal setup:
Using the provided VeReMi and CICIDS2017 datasets, the intrusion detection system adopts ensemble learning classifiers to tackle issues linked to multiple classification models. The CICIDS2017 set presents an enormous dataset based on many attacks, including DDoS and PortScans, recorded for numerous days. The VeReMi set contains a few comma-separated Values(CSV) files encapsulating acts of assault by a specific aggressor. The organisation phase required eliminating any records with null values denoted as NaNs or Infinities —to ensure the actual credibility of the models formed. The VeReMi labels are then transformed from numerical codifications into substantial descriptions, such as "Constant Attack" or "Random Attack.". The CICIDS2017 CSV file deduces that some attack types are formed into unique invasion types with a broader category. The viably divided datasets training and validation data sets at an 80:20 ratio to grant every variety proportion in training modelling. Normalisation by scaling methods like MinMaxScaler ensures the balance of size magnitudes of integer inputs into the numerical system to ensure the model is incapacitated to give visual bias. A blend of models like RandomForest and XGBoost has side benefits with Logistic Regression as the second phase model. For attacks less regularly known, like Heartbleed in VeReMil, the count of cross-validation regression folds is lowered to CV = 4 to prevent overfitting against the undersized represents of points. Metrics including accuracy, precision, recall, F1 score, and more sophisticated measurements like RMSE (Root-mean-square deviation) were used to assess the model.
5 Results and Evaluation
In the suggested model, accuracy is the most significant quality indicator. This precision is used to evaluate the model’s performance in the intrusion detection system. The performance of the suggested system has been assessed using accuracy (ACC), detection rate (DR), precision, and recall metrics. True positive (TP), true negative (TN), false positive (FP), and false negative (FN), accordingly.
During anomaly identification, the training and testing phases utilise distinct datasets. However, as evidenced by the results, the accuracy percentage is exceptionally high and can be considered outstanding. In addition, the other performance metric exhibits encouraging results.
@@ 5 Results and Evaluation
CICIDS2017
estimators = [(’rf’, RandomForestClassifier(n_estimators= 26, max_depth= 50,
min_samples_split= 2,
min_samples_leaf= 1, criterion= ’entropy’)),
(’xgb’, xgb.XGBClassifier(eval_metric=’rmse’, use_label_encoder=False))]
clf = StackingClassifier(estimators=estimators,
final_estimator=LogisticRegression(solver=’lbfgs’, max_iter=800000))
@@
estimators = [(’rf’, RandomForestClassifier(n_estimators= 100, max_depth= 50,
min_samples_split= 10,
min_samples_leaf= 2, criterion= ’entropy’)),
(’xgb’, xgb.XGBClassifier(eval_metric=’rmse’, use_label_encoder=False))]
clf = StackingClassifier(estimators=estimators,
final_estimator=LogisticRegression(solver=’lbfgs’, max_iter=800000),
@@
Table 5: Estimators for Multiclass Approach for small attack size.
Attacks CICIDS2017
Heartbleed
Infiltration
estimators = [(’rf’, RandomForestClassifier(n_estimators= 1, max_depth= 5, random_state=None,
min_samples_split= 11, min_samples_leaf= 1, criterion= ’entropy’)),
(’xgb’, xgb.XGBClassifier(eval_metric=’rmse’, use_label_encoder=False))]
clf = StackingClassifier(cv=4, estimators=estimators, final_estimator=LogisticRegression
(solver=’lbfgs’, max_iter=800000), stack_method=’auto’,
passthrough=False, n_jobs=-1)
Table 6: Model performance evaluation table for Binary Classification Approach
Datasets Accuracy Precision Recall F1 RMSE r2 Score Time (hh:mm:ss)
CICIDS2017 99.92% 99.75% 99.87% 99.81% 0.027 0.99 00:27:37
VeReMi 99.68% 99.60% 99.31% 99.45% 0.05 0.98 00:07:48
Table 7: Model performance evaluation using Multiclass Approach for CICIDS2017 dataset.
Attacks Accuracy Precision Recall F1 FN FP TN TP RMSE R2 Time
DoS 99.984 99.857 99.988 99.922 6 72 454193 50336 0.012 0.998 01:59:46
PortScan 99.958 99.386 99.987 99.686 4 196 454068 31757 0.020 0.993 01:24:27
DDoS 99.997 99.972 99.988 99.980 3 7 454257 25602 0.004 0.999 01:44:54
Brute Force 99.998 98.866 100 99.429 0 5 454259 436 0.003 0.988 01:31:02
Web Attack 99.998 98.863 99.770 99.315 1 5 454259 435 0.003 0.986 01:20:33
Bot 99.961 70.185 96.930 81.417 12 161 454104 379 0.019 0.557 01:37:20
Infiltration 100 100 100 100 0 0 454265 7 0 1 00:57:12
Heartbleed 99.999 80 100 88.888 0 1 908528 4 0.001 0.749 00:16:59
@@
Attacks Accuracy Precision Recall F1 FN FP TN TP RMSE R2 Time
Eventual Stop 99.705 99.355 98.065 98.706 219 72 87414 11100 0.054 0.970 00:09:42
Random 99.994 100 99.918 99.959 5 0 87486 6097 0.007 0.999 00:09:23
Constant Offset 99.989 100 99.835 99.917 10 0 87486 6085 0.010 0.998 00:08:21
Constant 100 100 100 100 0 0 87486 6095 0 1 00:02:10
Random Offset 99.962 99.897 99.507 99.702 29 6 87480 5863 0.019 0.993 00:13:32
Root-mean-square error (RMSE) is a regularly utilised measure of the changes between the expected values (sample or population values) and the actual values. The RMSE is the square root of the second instance of the deviations between expected and actual values or the quadratic mean of these deviations. When these deviations are computed across the test dataset used for the assessment, they are known as residuals; when estimated out-of-sample, they are known as errors (or prediction errors). The RMSE integrates the magnitudes of errors in forecasts for different data points into a familiar figure of projecting skill. Due to its scale dependence, RMSE is a measure of accuracy used to assess the prediction mistakes of various models for a single dataset and not between datasets. RMSE is always positive, and a value of 0 indicates that the model fits the data perfectly (which is practically never attained in practice). A smaller RMSE is often preferred over a greater one. However, because the measure depends on the scale of the employed numbers, comparisons between distinct data kinds are erroneous. The effect of each error on RMSE is equal to the error’s squared size; thus, more significant errors have a disproportionate effect on RMSE.
  • RMSE is the common deviation of the residuals.
  • RMSE shows average model prediction fault.
  • The lower values specify a good fit.
  • It is measured in the same units as the Target variable.
The percentage of the variation in the dependent variable that can be predicted from the independent variable is known as the coefficient of determination in statistics. It is symbolised by the symbols $R2$ or $r2$. It is a statistic applied to statistical models whose main objective is to either predict future outcomes or test hypotheses using data from other relevant sources. Based on the percentage of the overall variation in outcomes that the model is responsible for explaining, it gauges how effectively the model mirrors observed results. In most cases, the coefficient of determination lies between 0 and 1.
@@ 5.1 Experimental results
CICIDS2017 and VeReMi datasets are used to evaluate the performance of binary classification models and multiclass classification models. The binary classification method achieved good accuracy, precision, recall, and F1 score for both datasets. The model scored 99.92% accuracy, 99.75% precision, 99.87% recall, and 99.81% F1-score for CICIDS2017. The model scored 99.68% accuracy, 99.60% precision, 99.31% recall, and 99.45% F1-score for VeReMi. The RMSE and r2 values show that the model’s predictions were correct and fit the data well. The model trained on CICIDS2017 took longer than the model trained on VeReMi.
The multiclass classification strategy has been implemented using the CICIDS2017 and VeReMi datasets. The models were taught to recognise eight different types of attacks. For CICIDS2017, the model achieved high accuracy, precision, recall, and F1 score on all attacks. The model’s overall accuracy was 99.99%, and the lowest accuracy for an attack was 99.96% for PortScan. The model for VeReMi achieved great accuracy, precision, recall, and F1 score for all five types of attacks. The model’s overall accuracy of the model was 99.99%, and the lowest accuracy achieved for an attack was 99.71% for Eventual Stop.
The RMSE and r2 values indicate that the model’s predictions were correct and fit the data well. On the other hand, the model trained on the Bot attack for CICIDS2017 had a low r2 score of 0.557, indicating that the model’s predictions were not well-fitted to the data and the models’ training time varied according to the number of attacks to be classified. The CICIDS2017 model took longer, from 00:57:12 to 01:59:46, while the VeReMi model took between 00:02:10 and 00:13:32.
FN and FP are critical indicators of a classification model’s success since they accurately identify positive and negative situations. A successful model should have low FN and FP rates, indicating that it can detect positive and negative cases correctly. This model achieved this with an excellent score for multiclass classifications. To be more specific, this model performed effectively even though infiltration and heartbleed attack entries in CICIDS2017 were relatively shallow. The estimator table demonstrates how parameters were applied to produce the best results.
(a) Classification Report for CICIDS2017
(b) Classification Report for VeReMi
Figure 4: Classification Report by attacks.
These slight variances in performance for attack detection in the IDS model, using the CICIDS2017 and VeReMi datasets, can be enlightened better based on the model’s precision, recall, F1 scores, false positives, and false negatives. In the CICIDS2017 dataset, precision and F1 scores are significantly lower than for other attacks such as DoS and DDoS. This lower performance indicator might suggest that Bot attacks display patterns so akin to behaviour that the model finds it challenging to detect them without a false hit on benign activities as an attack. Therefore, the high number of positives for Bot attacks indicates that the model is more likely to misclassify traffic as malicious due to similarities to Bot-like behaviours. However, even though Heartbleed has the highest true positives and accuracy, the precision performance indicator is second-to-last in the dataset. The reduced precision of this particular attack indicates a Heartbleed in the dataset, and the model’s confidence is undermined by its experience with such examples. Alternatively, features of Heartbleed attacks may overlap significantly with behaviours, posing classification challenges. As highlighted in the dataset, except for the Eventual Stop attack, the detection performance is crucial for attacks, presenting a slightly lower precision and higher false negatives, implying that it is challenging for the model to detect this attack pattern consistently, perhaps because the specific characteristics of the Eventual Stop attack are not as straightforward or consistent. Besides, the high precision score for Random and Constant Offset attacks indicates that these attacks have consistent patterns, which the model can learn quickly. These might have features that differ significantly from behaviour. Hence, they are easily identified by the model. Moreover, the false negative rates will indicate a balance between specificity and sensitivity in the model. High false negative rates, as observed in the Eventual Stop attack, suggest that the model is precise in its predictions but is not sensitive enough to detect this attack in some cases. Random Forest and XGBoost as base learners and Logistic Regression as a learner have shown model effectiveness, considering now the performance evaluation for CICIDS2017 and VeReMi datasets. The model implementation involved a combination of these algorithms purposely to use their strengths and mitigate their weaknesses through stacking. The variance in model effectiveness against types of attacks is influenced by the complexity and frequency of attack patterns, the degree of similarity among harmful and non-harmful features, and the data volume and quality that have been used to train the model. These factors contribute to how efficiently the model has learned, ultimately influencing its accuracy in detecting and classifying types of attacks.
@@
In the end, in the CICIDS2017 and VeReMi datasets, binary and multiclass classification models performed well in identifying attacks. The models’ high accuracy, precision, recall, and F1-score and the low RMSE and high r2 values show that the model’s predictions were accurate and well-fitted for the data. The training duration varies according to the size of the dataset and the number of attacks to be classified.
@@
Moreover, to further solidify the claim that the proposed approach was robust and efficient, we compared our stacked ensemble model with other methods, as presented in Table 3. For instance, Gad et al. (2021) attained 99.1% accuracy while employing XGBoost after feature selection, Türkoğlu et al. (2022) obtained 99.35% accuracy with Decision Tree after MRMR feature selection, Zeng et al. (2019) used DeepVCM with CNN and LSTM to realise 98.5% accuracy, and Aneja et al. (2018) developed an ANN with a genetic algorithm to get 99% accuracy. Comparatively, our proposed stacked ensemble model introduced higher accuracy levels at 99.92% with the CICIDS2017 dataset and 99.68% with the VeReMi dataset.
5.2 Challenges and Experiences to Implementation
Depending on database size and hardware capacity, the elapsed time is a big challenge in the classification phase. Hardware capacity can be upgraded to get quick results, but it will not be cost-effective for customers. The two publicly available datasets, VeReMi and CICIDS2017, are the primary source of our study’s glaring errors. Although these synthetic datasets are widely used in research, they cannot fully represent the intricacy of actual vehicle networks. More precisely, other notable attack types, such as denial of service or time-sensitive threats, are not included in the VeReMi dataset. It primarily focuses on misbehaviour attacks, such as fabricating location data. Due to the constraint mentioned above, our model may be unable to identify as many harmful activities as possible in real-world scenarios.
Developing a stacked ensemble is a proposed learning method that uses Random Forest, XGBoost, and Logistic Regression algorithms to obtain a reliable Intrusion Detection System for Software-Defined VANETs. Random Forest and XGBoost are computationally expensive during training because they have many trees and extensive data sorting. In comparison, Logistic Regression is simple but requires more iterations to converge. The storage requirement pertains to the number of trees and storage units and their weight matrices in Logistic Regression. The complexity of the ensemble model will affect the individual classifiers’ training and storage space. The model is efficient and effective as an IDS. The approach is practical in all aspects. It has enhanced the various trade-offs in achieving high accuracy and higher performance in terms of efficiency, rendering it appropriate for the operation of the dynamic and resource-constrained network, which are the VANETs. Generally, from the findings, VANETs have various requirements and windows for realising the effectiveness of improving an IDS.
5.3 General Implementation method in SDN-based VANET
The SDN controller has two unique interfaces, one on the northbound side that communicates with the cloud, orchestration, and management layer and the other on the southbound side that talks to the data plane layer, as shown in Figure 1. In most of the SDN controllers that are now available, the northbound interface is standardised as REST APIs. Only the vendor’s implementation of the southbound interface is supported. Some use OpenFlow, BGP, Netconf, and other open standards-based protocols, while others use proprietary interfaces. An SDN controller is a software scheme or group of systems that collectively offers the following:
(A) A database will be used to manage the state of the network and, in some situations, to distribute this state. These databases act as a store for data obtained by the managed network components and associated software, as well as data managed by SDN applications (including network state, learned topology, specific ephemeral arrangement data, and control session data). In some circumstances, the controller could have several targeted data procedures. Different in-memory database techniques may also be used in other situations.
(B) A large data type depicts the connections between controlled resources, rules, and other services the controller offers.
(C) An application can access the controller services through a modern, frequently RESTful (Representational State Transfer) application programming interface (API). This makes the vast majority of controller-to-application interactions more accessible. The database schema representing the controller’s services and features should be used to construct this interface. In certain situations, an application framework that generates API code from a model contains both the controller and its API. Some systems offer sophisticated DevOps (software development and operations) that permit the development of core capabilities and the publishing of APIs for new modules, including those that facilitate the continuous expansion of controller abilities.
(D)Between the controller and the connected agents in the network elements, a secure TCP control session
(E) A protocol that adheres to standards for the delivery of application-driven network state on network components
(F) A path computation system, a device, topology, service discovery method, and possibly additional network- or resource-centric information services.
The baseline is this IDS solution can be installed in the OpenFlow switch in the control plane layer.
Advantages and Limitations of the proposed method
Advantages:
  • Both models are a good option when a single classifier cannot predict with optimum accuracy.
  • Three-stage ensemble method will reduce the risk of Inference Attacks. Salem et al. (2018)
  • Quick run-time with impressive accuracy.
  • Even performs well against small-scale attacks.
Limitation:
It cannot detect encrypted and unknown attacks.

### v5

@@ 4.1 Supervised ML Approaches
The Random Forest Classifier is an effective machine learning algorithm that is part of the ensemble learning category. It works by creating numerous decision trees during the training process and combining their predictions to enhance overall accuracy. Each tree is built using a random sample of the data, which helps to introduce variety and minimize the chances of overfitting. The final classification result is determined by a majority vote, where the most frequently predicted class among all the trees is chosen. This method not only improves accuracy but also adds resilience against noise and fluctuations in the dataset. Random Forest classifiers are popular across various fields, such as finance, healthcare, and cybersecurity, because they can manage large datasets with many features and provide insights into the importance of different features, making them both interpretable and effective for tasks involving classification and regression.
@@
LR is a linear model used to classify instead of predict. A logistic function is used for probabilities that describe the possible results of a single trial in this model. Typically, a sigmoid curve with the shape of a Sigmoid curve is called a logistic function or logistic curve (Wikipedia (2023)).
CatBoost:
As noted by Dorogush et al. (2018), CatBoost is a gradient-boosting algorithm developed by Yandex that excels in handling categorical features, making it particularly suitable for datasets with heterogeneous data types. Unlike traditional boosting methods, CatBoost employs an innovative approach called ordered boosting, which mitigates overfitting by utilizing a permutation-based method for training. Its efficiency is further enhanced through support for both numerical and categorical data, eliminating the need for extensive preprocessing. With state-of-the-art performance across various machine learning tasks, CatBoost has gained popularity in both academia and industry, offering robustness and interpretability in model training and evaluation, particularly in the context of complex datasets.
In essence, Random Forest is particularly strong at managing high-dimensional data and is known for its resilience against overfitting. On the other hand, CatBoost shines when dealing with categorical features and effectively reduces gradient bias. By stacking these two models, we aim to improve both detection performance and interpretability within VANETs.
5 Methodology
@@ 5.1 Technical Implementation Methodology
This study addresses multiclass classification problems using two approaches: binary decision reduction and multiclass methodology. The performances of these two approaches will be discussed below.
5.2 Data Processing
5.2.1 Binary Decision Reduction
Before inputting the data into machine learning algorithms, it is essential to perform cleaning and preparation steps to ensure high performance. Unnecessary features may detract from the chosen ML method’s effectiveness. We handled missing values, which are common in datasets, by replacing NaNs and infinities with appropriate substitutes. After grouping similar attacks and splitting the dataset into training and testing parts, we employed a binary classifier to label the dependent columns as either BENIGN or ATTACK. LabelEncoder normalized the labels, ensuring that the dataset contained values only between 0 and $n_{\text{classes}}-1$. Additionally, MinMax Scaling was applied to mitigate the effects of variables measured on different scales, helping to prevent bias.
5.2.2 General Multiclass Methodology
@@
Distinct datasets were constructed by isolating specific attacks while discarding the remainder, thereby creating unique data combinations for each attack versus BENIGN data.
5.3 Feature Selection
@@
The SelectKBest class from the scikit-learn library employs various statistical tests to identify specific features. In this experiment, the chi-squared ($\chi^{2}$) statistical test was utilized to select the ten best features based on non-negative characteristics. The Chi-Square test quantifies the disparity between observed and expected data, allowing us to assess whether the discrepancies between categorical variables result from chance or an underlying association.
5.4 Hyperparameter Optimization: Bayesian Optimization with Gaussian Process
BayesSearchCV integrates “fit” and “scoring” methods, optimizing parameters through a cross-validated search of the parameter settings. Unlike GridSearchCV, which tests all parameter values, this approach experiments with several parameter settings drawn from selected distributions. The number of iterations ($n_{\text{iter}}$) determines how many parameter settings will be assessed, leading to a more efficient exploration of the hyperparameter space.
The hyperparameters can be continuous, discrete, or categorical. In this study, we implemented hyperparameter optimization for the stacking base classifier.
5.5 Ensemble Stacking Classifier
Ensemble learning encompasses three primary methods: bagging, boosting, and stacking. Bagging and boosting utilize alternative voting methods, while stacking serves as a versatile framework, integrating multiple ensemble approaches. This study utilized a stacking approach with Random Forest and CatBoost as base learners and Logistic Regression as the final estimator. The configuration for the stacking classifier is as Table 3:
@@ 5.5 Ensemble Stacking Classifier
estimators = {(’rf’, RandomForestClassifier(
n_estimators=50, max_depth=20,
min_samples_split=10, min_samples_leaf=2,
criterion=’entropy’, n_jobs=-1)),
(’catboost’, CatBoostClassifier(
iterations=100, learning_rate=0.1,
depth=6, verbose=0)}
clf = StackingClassifier(
cv=3, estimators=estimators,
final_estimator=LogisticRegression(
solver=’lbfgs’, max_iter=1000),
@@
The choice of base learners is crucial for the stacking process, highlighting the importance of employing multiple models rather than relying on a single one. The stacking method uniquely employs meta-level learning for the final decision-making process.
5.6 Experimental Setup
With the help of the VeReMi datasets, our intrusion detection system employs ensemble learning classifiers to tackle multiclass classification challenges.
The VeReMi dataset comprises multiple CSV files capturing different types of attacks. Data preprocessing involved eliminating records with null values (NaNs or Infinities) to ensure model reliability. Labels were transformed from numerical codes to meaningful descriptions (e.g., “Constant Attack”, “Random Attack”). The dataset was divided into training and validation sets in an 80:20 ratio to ensure a balanced representation of attack types. Normalization using MinMaxScaler ensured that all features were scaled appropriately, thereby eliminating visual biases in the model’s predictions. Our approach involves combining models like RandomForest and CatBoost, with Logistic Regression serving as the meta-classifier. To address underrepresented attacks, cross-validation folds were set to CV = 3, helping to prevent overfitting against smaller sample sizes.
6 Results and Evaluation
In our proposed model, accuracy serves as the primary metric for assessing performance within the Intrusion Detection System (IDS). The evaluation metrics include accuracy (ACC), detection rate (DR), precision, and recall, based on true positive (TP), true negative (TN), false positive (FP), and false negative (FN) counts.
The performance evaluation leverages distinct datasets for the training and testing phases of anomaly detection. As evidenced by the results, the achieved accuracy percentage is notably high, warranting recognition as outstanding. Additionally, the other performance metrics exhibit similarly encouraging results.
Table 4: Performance evaluation for Binary Classification
Datasets Accuracy Precision Recall F1
VeReMi $99.68\%$ $99.60\%$ $99.31\%$ $99.45\%$
@@ 6 Results and Evaluation
Attacks Accuracy Precision (%) Recall (%) F1 (%)
BENIGN $97.65\%$ $100\%$ $98\%$ $99\%$
Constant Attack $100\%$ $100\%$ $100\%$ $100\%$
Constant Offset Attack $99.74\%$ $99\%$ $100\%$ $99\%$
Eventual Stop Attack $97.14\%$ $78\%$ $97\%$ $87\%$
Random Attack $99.95\%$ $100\%$ $100\%$ $100\%$
Random Offset Attack $99.10\%$ $98\%$ $99\%$ $99\%$
Accuracy $98.11\%$
Macro Avg $97.64\%$ $96.17\%$ $99.57\%$ $97.14\%$
Weighted Avg $98.11\%$ $98.29\%$ $98.38\%$ $98.28\%$
@@ 6.1 Experimental Results
The evaluation of the VeReMi dataset provides comprehensive insights into the performance of both binary and multiclass classification models. The binary classification approach achieved remarkable metrics, with accuracy reaching $99.68\%$, precision at $99.60\%$, recall at $99.31\%$, and an F1 score of $99.45\%$ for the VeReMi dataset. In contrast, the multiclass classification strategy adeptly identified various attack types, with the model achieving an overall accuracy of $98.11\%$. The lowest accuracy observed was $97.14\%$ for the Eventual Stop Attack, highlighting an area for potential improvement.
False negatives (FN) and false positives (FP) are critical indicators of a classification model’s success, as they represent the model’s ability to accurately identify both positive and negative instances. An effective model should maintain low FN and FP rates, reflecting its capacity to correctly detect positive and negative cases. Our model demonstrated exceptional performance, particularly in multiclass classifications. The estimator table outlines the parameters applied to achieve these impressive results.
Figure 4: Confusion Matrix by Attacks for Multiclass Classification.
The slight performance variances observed in attack detection within the IDS model using the VeReMi dataset can be further elucidated through metrics such as precision, recall, F1 scores, false positives, and false negatives. As indicated in the dataset, the model exhibited slightly lower precision and higher false negatives for the Eventual Stop Attack, suggesting inherent challenges in consistently detecting this specific attack pattern. This difficulty may stem from the less straightforward characteristics associated with the Eventual Stop Attack. Conversely, the high precision scores for the Random and Constant Offset attacks suggest that these attack patterns present consistent features, enabling the model to learn them efficiently.
Moreover, false negative rates indicate a balance between specificity and sensitivity within the model. The elevated false negative rate for the Eventual Stop Attack implies that while the model may be precise in its predictions, it lacks the sensitivity necessary to detect this attack consistently. The use of Random Forest and CatBoost as base learners, complemented by Logistic Regression as a final estimator, has enhanced model effectiveness, particularly in evaluating performance against the VeReMi dataset. The strategic implementation of these algorithms allows for the leveraging of their strengths while mitigating their weaknesses through a stacking approach. The variance in model effectiveness across different attack types is influenced by factors such as the complexity and frequency of attack patterns, the degree of similarity among harmful and benign features, as well as the volume and quality of data utilized for training. Collectively, these factors play a crucial role in determining how effectively the model learns, ultimately shaping its accuracy in detecting and classifying various attack types.
@@
In conclusion, both binary and multiclass classification models performed admirably in detecting attacks within the VeReMi dataset. The high accuracy, precision, recall, and F1 scores, alongside low RMSE and high $R^{2}$ values, affirm that the model’s predictions were accurate and well-fitted to the data. The training duration varied according to the dataset size and the number of attack types classified.
@@
To substantiate the robustness and efficiency of our approach, we conducted comparative analyses against other established methods, detailed in Table 3. For instance, Gad et al. (2021) achieved an accuracy of $99.1\%$ using XGBoost after feature selection; Türkoğlu et al. (2022) attained $99.35\%$ accuracy with a Decision Tree following MRMR feature selection; Zeng et al. (2019b) realized $98.5\%$ accuracy with a combination of CNN and LSTM in DeepVCM; and Aneja et al. (2018) developed an ANN with a genetic algorithm, achieving $99\%$ accuracy. In contrast, our proposed stacked ensemble model surpassed these benchmarks, achieving an impressive accuracy of $99.68\%$ with the VeReMi dataset.
(a) SHAP Analysis Plot for CatBoost
(b) SHAP Analysis Plot for Random Forest
Figure 5: Feature Impact and Misclassification Analysis for Multiclass Classification.
6.2 Feature Impact and Misclassification
The confusion matrix reveals the model’s strong capability in classifying network behaviours, particularly in accurately identifying benign traffic, with 70,834 correct classifications. However, there are instances where benign traffic is misclassified as Eventual Stop Attacks (1,549 cases), highlighting areas that need enhancement. On the other hand, categories such as Constant Attack, Constant Offset Attack, and Random Attack demonstrate remarkable accuracy, with very few misclassifications. Although the model correctly identifies the Eventual Stop Attack 5,601 times, it occasionally misclassifies this category as benign, signalling the necessity for improved detection strategies.
This experiment used SHAP analysis to assess how well the CatBoostClassifier and RandomForestClassifier models could distinguish between benign and attack traffic. The SHAP summary plots revealed important insights, showing which features impacted the model’s decisions most. For CatBoost, features like pos-y1, pos-x1, and spd-y1 stood out, especially when their values suggested abnormal activity. Similarly, RandomForest heavily relied on features like sender_1, messageID, and spd-x1, where overlapping patterns between benign and attack data pointed to possible misclassification issues.
SHAP interaction plots allowed for further examination of feature independence and vulnerability. A notable interaction was found between pos-y1 and spd-x1, suggesting a dependency between these features. This raises a critical issue, as altering one feature could affect the other, leading to incorrect classifications. This is an important consideration, as attackers could potentially exploit these dependencies to deceive the model.
Additionally, the overlapping SHAP values between benign and attack data clarified why the model sometimes misclassified attack traffic as benign. These similarities in feature distributions between the two classes complicated the classification process. This suggests that using a two-step classification system—first identifying traffic as either benign or attack, and then classifying specific attack types—could help reduce the risk of critical errors, such as attacks being mistakenly identified as benign, which is a significant concern in cybersecurity.
The SHAP analyses also helped explain the misclassifications, as the similar feature importance between the classes indicated that the model’s errors were understandable. This highlights the need for improved feature selection or refining the decision boundaries, particularly for features where overlap occurs, to enhance the model’s ability to accurately detect attacks.
7 Limitations and Future Scope
7.1 Expanded Limitations
One of the major hurdles we encountered in this study stems from our reliance on SHAP analysis for understanding how our model makes decisions. While SHAP is a powerful tool for identifying the significance of different features, it doesn’t fully capture the complex relationships and dependencies that exist among them. For example, features like position coordinates (pos-y1) and speed (spd-x1) may show overlapping SHAP values, which can lead to confusion in classification tasks. This overlap can create ambiguity that misguides the model in critical situations, rendering it vulnerable to adversarial attacks that exploit these hidden connections.
7.2 Future Directions
As we look ahead, we see tremendous potential for our framework to evolve by incorporating encrypted data or data obfuscation. This move would not only enhance privacy but also ensure that threat detection remains robust. Furthermore, putting our model to the test on real-time data from VANETs could provide invaluable insights into its scalability and reliability in high-stakes, real-world scenarios. We also believe it is essential to explore alternative methods of explainability that prioritize feature independence. By addressing some of the shortcomings associated with SHAP analysis, we could better equip our model to handle more complex challenges. Ultimately, our goal is to build a more resilient security framework that meets the demands of intelligent transportation systems, paving the way for safer and smarter roadways.